OpenStudy (anonymous):

Solve the equation: 2sinx-1=0

OpenStudy (anonymous):

2sin(x-1) or 2sin(x) - 1?

OpenStudy (anonymous):

sin(x) = 1/2 x = ?

OpenStudy (anonymous):

so what's the pi

OpenStudy (anonymous):

What's your answer? Does it have pi in it?

OpenStudy (anonymous):

theses are my opions pi/3, 4pi/5 pi/6, 5pi/62pi/5, 3pi/4 I dont know which one

OpenStudy (anonymous):

If you put it into a calculator, you get 30 degrees. Drawing a circle and putting pi/6,/pi/4, pi/3 on the circle for 30 degrees, 45 degrees and 60 degrees, respectively will give you a good picture of where those line up with values for the sin function. I hope that made sense. So, 30 degrees = pi/6

OpenStudy (anonymous):

it's \[\pi/6 \] or \[5\pi/6\]

OpenStudy (fellowroot):